find angle A of the oblique triangle with sides a=13, b=14, and c=15
Answer by mducky2(62)

If you need to find an angle when three sides of an oblique triangle are known, use the law of cosines. The equation is:
a2 = b2 + c2 - 2bc cos A

Solve for A
2bc cos A = b2 + c2 - a2
cos A = (b2 + c2 - a2)/(2bc)
A = cos-1[(b2 + c2 - a2)/(2bc)]
A = cos-1[(142 + 152 - 132)/(2*14*15)]
A = 53.13°